OpenStudy (anonymous):

What is the value of 16 1/2? Answer a. 4 b. 1/8 c. 8 d. 1/4

OpenStudy (anonymous):

is the question 16 * 1/2 ?

OpenStudy (anonymous):

a.4\[\sqrt{16}\]

OpenStudy (anonymous):

I think he means \[16^{\frac{ 1 }{ 2 }}\] which is \[\sqrt{16}\] So the answer is 4.